# Quantum reflection time and the Goos-Hänchen effect

### v1

@@ Appendix B. Gaussian wave packets
${\langle}x|\Psi(0){\rangle}=\left(\frac{\pi\Delta x^{2}}{2}\right)^{-1/4}\exp\left[-\frac{(x-x_{0})^{2}}{\Delta x^{2}}+ip(x-x_{0})\right]\quad$ (67)
where $\Delta x=2/\Delta k$. For $p<\Delta k$, the initial WP contains also negative momenta, even though its move towards the barrier. Similarly, with $p<0$, the WP moving now away from the barrier contains positive momenta. The first term in Eq.(62), describing free motion of the initial wave packet, is now given by
@@
$G_{0}(x-\frac{pt}{m},t)=\left(\frac{2\Delta x^{2}}{\pi\Delta x_{t}^{4}}\right)^{1/4}\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x_{t}^{2}}\right],\quad$ (69)
$\Delta x_{t}\equiv\sqrt{\Delta x^{2}+\frac{2it}{m}},\quad\quad\quad\quad\quad\quad\quad$
@@
$G_{0}(x-\frac{pt}{m},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x^{2}}\right].\quad$ (70)
Equations (68)-(70) are valid, we repeat if the initial Gaussian state (66) contains momenta in the direction opposite to the WP’s velocity $v$.
@@ Appendix C. Free motion and reflection from an infinite (hard) wall
More interesting is the case of reflection by an infinite wall placed at some $0<a$, $V(x)=0$ for $x<a$, and $V(x)=\infty$ for $x\geq a$. Now we have
@@
$\xi(x^{\prime})=-\delta(x^{\prime}-2a),\quad\quad\quad\quad\quad\quad\quad$
@@
$\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dkR_{l}(k){\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}=$ (72)
$-\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{-ik(x-2a)-iE(k)t}.$
@@ Appendix D. Quantum measurements
$\hat{H}^{P}=\hat{\lambda}^{2}/2\mu,\quad\hat{H}_{int}=\hat{\lambda}\hat{B}\delta(t+0),$ (74)
where $\mu$ is the pointer’s mass, and $\hat{B}=\sum_{i}|b_{i}{\rangle}B_{i}{\langle}b_{i}|$ is the system’s operator to be measured. The system is pre- and post-selected in states $|I{\rangle}$ and $|F{\rangle}$ at $t=0$ and $t$, respectively. At $t=0$ the pointer is prepared in the Gaussian state (67) $x\to f$, $p\to{\lambda}$, $\Delta x\to\Delta f$, $x_{0}\to f_{0}=0$]. At time $t$ the observer accurately determines the pointer’s position, hoping in this way to learn something about the value of $\hat{B}$ at the time the system interacted with the pointer. The pointer’s state after the system was found in $|F{\rangle}$ at $t$, is
${\langle}F|\Phi(t){\rangle}=\sum_{i}\Psi(f-B_{i},t)\mathcal{A}^{S}(F\leftarrow b_{i}\leftarrow I),$ (75)
where $\Psi(f-B_{i},t)$ is given by Eqs.(68)-(69), and $\mathcal{A}_{i}^{S}$ is the system’s transition amplitude,
@@
where $\xi(f^{\prime})$ is the probability amplitude of the quantity, represented by the operator $\hat{B}$, having a value $f^{\prime}$. Suppose we can neglect the spreading of the pointer’s wave packet, prepared with mean momentum ${\lambda}=0$, $\Psi(f-f^{\prime},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(f-f^{\prime})^{2}}{\Delta f^{2}}\right]\equiv G_{0}(f-f^{\prime})$. Now only the values in the interval $f-\Delta f\lesssim f^{\prime}\lesssim f+\Delta f$ contribute to $\Phi(f,t)$ and by determining $f$ accurately one measures $\hat{B}$ to a (quantum) accuracy $\Delta f$. Note that as $\Delta f\to\infty$ the probability of the post-selection, $P(F)$, is no longer affected by the pointer, $P(F)\to|{\langle}F|\exp(-i\hat{H}^{S}t|I{\rangle}|^{2}$.
@@
and finding the pointer at $f$ one can still know that $\hat{B}$ had the value roughly in the interval $(f-\Delta f,f+\Delta f)$. However, now sending $\Delta f\to\infty$ does not fully decouple the pointer from the system, as $P(F)\to|\int df^{\prime}\exp(-i{\lambda}f^{\prime})\xi(f^{\prime})|^{2}\neq|{\langle}F|\exp(-i\hat{H}^{S}t|I{\rangle}$. The similarity between Eqs.(78) and (8) provides the required analogy.
@@ Appendix E. Rectangular barrier and potential step
Consider a rectangular barrier $V(x)$ of a hight $V=const$, occupying the region $0\leq x\leq d$. For a particle of mass $m$, the reflection amplitude is well known to be
@@
where $q(k)\equiv\sqrt{k^{2}-2mV}\equiv\sqrt{k^{2}-W^{2}}$. The amplitude is an entire function in the complex $k$-plane, unchanged by replacement $q\to-q$ and satisfying, for real $k$ , $R(-k)=R^{*}(k)$. It has poles in the lower half plane [complex zeroes of the denominator in Eq.(80)], associated with the resonances at the energy above the barrier, $k^{2}/2m>V$. Thus to calculate $\xi(x^{\prime})=(2\pi)^{-1}\int_{-\infty}^{\infty}dk\exp(-ikx^{\prime})R(k)$ we may shift the contour of integration into the upper half-plane by putting $k=k+i\epsilon$. $\epsilon\to 0$, so $\xi(x^{\prime})=0$ for $x^{\prime}<0$. Note that on the contour shown in Fig.6 q(k)=q(k)q(-k)=q(k) for $|k|\leq W$, and q(k)=q(k)q(-k)=-q(k) for $|k|>W$,
@@
Figure 6: For $x^{\prime}>0$, the contour of integration running just above the real $k$-axis can be replaced by a loop around the branching point at $k=\pm\sqrt{2mV}$
@@
$\xi(x^{\prime}\geq 0)=(2\pi)^{-1}\oint dkR(k)\exp(-ikx^{\prime})\quad\quad$ (82)
$=-\frac{4W}{\pi}\int_{0}^{1}dtt\sqrt{1-t^{2}}\sin(Wx^{\prime}t),\quad\quad\quad\quad$
$=\frac{2}{\sqrt{2mV}}\frac{d}{dx^{\prime}}\left(\frac{J_{1}(\sqrt{2mV}x^{\prime})}{x^{\prime}}\right),\quad\quad\quad\quad$
and $\xi(x^{\prime}<0)\equiv 0$. In (82) $J_{1}(z)$ is the Bessel function of the first order Abram . It is easy to check that $\int_{0}^{\infty}\xi(x^{\prime})=R(p=0)=-1$. Fimally, using Eqs. 9.1.30 of Abram , one obtains
@@ Appendix F. Centre-of-mass of the reflected wave packet
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{\infty}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{\infty}dx\,|\Psi_{R}(x,t)|^{2}}$ (84)
@@
$F(k,t)\equiv R_{l}(k){\langle}k|\Psi(0){\rangle}\exp[-iE(k)t].$
@@ Appendix G. Gaussian beam in paraxial approximation
If the spread of the tranverse moment $\Delta q$ is small compared to $p$, for $x>x_{0}$ we have
@@
now accounts for beam’s broadening. Neglecting the broadening yields yields Eq.(41).

### v2

@@ Appendix B. Gaussian wave packets
${\langle}x|\Psi(0){\rangle}=\left(\frac{\pi\Delta x^{2}}{2}\right)^{-1/4}\exp\left[-\frac{(x-x_{0})^{2}}{\Delta x^{2}}+ip(x-x_{0})\right]\quad\quad$ (68)
where $\Delta x=2/\Delta k$. For $p<\Delta k$, the initial WP contains also negative momenta, even though it moves towards the barrier. Similarly, with $p<0$, the WP moving now away from the barrier contains positive momenta. The first term in Eq.(63), describing free motion of the initial wave packet, is now given by
@@
$G_{0}\left(x-\frac{pt}{m},t\right)=\left(\frac{2\Delta x^{2}}{\pi\Delta x_{t}^{4}}\right)^{1/4}$ $\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x_{t}^{2}}\right],$
$\Delta x_{t}\equiv\sqrt{\Delta x^{2}+\frac{2it}{m}},$
@@
$G_{0}(x-\frac{pt}{m},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x^{2}}\right].\quad\quad$ (71)
Equations (69)-(71) are valid, we repeat, even if the initial Gaussian state (67) contains momenta in the direction opposite to the WP’s velocity $v$.
@@ Appendix C. Free motion and reflection from an infinite (hard) wall
More interesting is the case of reflection by an infinite wall placed at some $a>0$, $V(x)=0$ for $x<a$, and $V(x)=\infty$ for $x\geq a$. Now we have
@@
$\xi(x^{\prime})$ $=-\delta(x^{\prime}-2a),$
@@
$\Psi_{R}(x,t)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dkR_{l}(k){\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}$ (73)
$=-\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{-ik(x-2a)-iE(k)t}.$
@@ Appendix D. Quantum measurements
$\hat{H}^{P}=\hat{\lambda}^{2}/2\mu,\quad\hat{H}_{int}=\hat{\lambda}\hat{B}\delta(t+\epsilon),\quad\epsilon\to 0,$ (75)
where $\mu$ is the pointer’s mass, and $\hat{B}=\sum_{i}|b_{i}{\rangle}B_{i}{\langle}b_{i}|$ is the system’s operator to be measured. The system is pre- and post-selected in states $|I{\rangle}$ and $|F{\rangle}$ at $t=0$ and $t$, respectively. At $t=0$ the pointer is prepared in the Gaussian state (55) $x\to f$, $p\to{\lambda}$, $\Delta x\to\Delta f$, $x_{0}\to f_{0}=0$], so that the composite’s initial state is given by $|\Phi(0){\rangle}=|I{\rangle}|\Psi_{0}{\rangle}$. At time $t$ the observer accurately determines the pointer’s position, hoping in this way to learn something about the value of $\hat{B}$ at the time the system interacted with the pointer. The pointer’s state after the system was found in $|F{\rangle}$ at $t$, is
${\langle}f|{\langle}F|\Phi(t){\rangle}=\sum_{i}\Psi(f-B_{i},t)\mathcal{A}^{S}(F\leftarrow b_{i}\leftarrow I),$ (76)
where $\Psi(f-B_{i},t)$ is given by Eqs.(69)-(Appendix B. Gaussian wave packets), and $\mathcal{A}_{i}^{S}$ is the system’s transition amplitude,
@@
where $\xi(f^{\prime})$ is the probability amplitude of the quantity, represented by the operator $\hat{B}$, having a value $f^{\prime}$. Suppose we can neglect the spreading of the pointer’s wave packet, prepared with mean momentum ${\lambda}=0$, $\Psi(f-f^{\prime},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(f-f^{\prime})^{2}}{\Delta f^{2}}\right]\equiv G_{0}(f-f^{\prime})$. Now only the values in the interval $f-\Delta f\lesssim f^{\prime}\lesssim f+\Delta f$ contribute to $\Phi(f,t)$ and by determining $f$ accurately one measures $\hat{B}$ to a (quantum) accuracy $\Delta f$. Note that as $\Delta f\to\infty$ the probability of the post-selection, $P(F)$, is no longer affected by the pointer, $P(F)\to|{\langle}F|\exp(-i\hat{H}^{S}t)|I{\rangle}|^{2}$.
@@
and finding the pointer at $f$ one can still know that $\hat{B}$ had the value roughly in the interval $(f-\Delta f,f+\Delta f)$. However, now sending $\Delta f\to\infty$ does not fully decouple the pointer from the system, as $P(F)\to|\int df^{\prime}\exp(-i{\lambda}f^{\prime})\xi(f^{\prime})|^{2}\neq|{\langle}F|\exp(-i\hat{H}^{S}t)|I{\rangle}$. The similarity between Eqs.(79) and (8) provides the required analogy.
@@ Appendix E. Rectangular barrier and potential step
Consider a rectangular barrier $V(x)$ of height $V=const$, occupying the region $0\leq x\leq d$. For a particle of mass $m$, the reflection amplitude is well known to be
@@
where $q(k)\equiv\sqrt{k^{2}-2mV}\equiv\sqrt{k^{2}-W^{2}}$. The amplitude is an entire function in the complex $k$-plane, unchanged by replacement $q\to-q$ and satisfying, for real $k$ , $R(-k)=R^{*}(k)$. For finite $d$, the reflection coefficient has zeros in the real $k$ axis and isolated poles in the lower half plane [complex zeroes of the denominator in Eq.(81)]. The zeros at $k_{l}=\pm\sqrt{(l\pi/d)^{2}+W^{2}}$ correspond to energies at which the barrier is perfectly transparent. The poles, corresponding to the zeros of the denominator in Eq.(81), are associated with the resonances (or quasi-stationary states) BZP at the energies above the barrier, $k^{2}/2m>V$.
Figure 6: The exact (numerical) poles of Eq.(81) evaluated for $d=10,15,20,30,50$ (circles) in orange, magenta, green, blue and red respectively. Also shown the zeros of $R(k)$ for $d=50$ (red crosses). The dashed grey lines show the approximations from Eq.(82).
For large $d$, their approximate position in the $k$-plane is given by,
$k_{n}$ $\approx\chi_{n}(d)-\frac{1}{d}\thinspace\frac{\beta_{n}(d)}{\chi_{n}(d)}\times\bigg(i\log\left|{\frac{\chi_{n}(d)+\beta_{n}(d)}{\chi_{n}(d)-\beta_{n}(d)}}\right|$ (82)
$-\arg\left(\frac{\chi_{n}(d)+\beta_{n}(d)}{\chi_{n}(d)-\beta_{n}(d)}\right)\bigg)+\mathcal{O}(d^{-2}),$
where $\beta_{n}(d)\equiv n\pi/d$ and $\chi_{n}(d)\equiv\sqrt{\beta_{n}(d)+W^{2}}$ (see Fig.6). As $d\to\infty$ the poles approach the real axis and together with the zeros form cuts drawn from the branch points $k=\pm\sqrt{W}$ to $\pm\infty$. A similar behaviour of densely coalescing poles was reported, e.g. in Dense_poles . This branch cut corresponds to the one present in the reflection amplitude for the step potential found in Eq.(83). Thus, to calculate $\xi(x^{\prime})=(2\pi)^{-1}\int_{-\infty}^{\infty}dk\exp(-ikx^{\prime})R(k)$ we may shift the contour of integration into the upper half-plane by putting $k=k+i\epsilon$. $\epsilon\to 0$, so $\xi(x^{\prime})=0$ for $x^{\prime}<0$. Note that on the contour shown in Fig.7, q(k)=q(k)q(-k)=q(k) for $|k|\leq W$, and q(k)=q(k)q(-k)=-q(k) for $|k|>W$.
@@
Figure 7: For $x^{\prime}>0$, the contour of integration running just above the real $k$-axis can be replaced by a loop around the branching points at $k=\pm\sqrt{2mV}$
@@
$\xi(x^{\prime}\geq 0)$ $=(2\pi)^{-1}\oint dkR(k)\exp(-ikx^{\prime})$ (84)
$=-\frac{4W}{\pi}\int_{0}^{1}dtt\sqrt{1-t^{2}}\sin(Wx^{\prime}t),$
$=\frac{2}{\sqrt{2mV}}\frac{d}{dx^{\prime}}\left(\frac{J_{1}(\sqrt{2mV}x^{\prime})}{x^{\prime}}\right),$
and $\xi(x^{\prime}<0)\equiv 0$. In (84) $J_{1}(z)$ is the Bessel function of the first order Abram . It is easy to check that $\int_{0}^{\infty}\xi(x^{\prime})=R(p=0)=-1$. Finally, using Eqs. 9.1.30 of Abram , one obtains
@@ Appendix F. Centre-of-mass of the reflected wave packet
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{\infty}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{\infty}dx\,|\Psi_{R}(x,t)|^{2}},$ (86)
@@
$F(k,t)\equiv R_{l}(k){\langle}k|\Psi(0){\rangle}\exp[-iE(k)t]$
@@ Appendix G. Gaussian beam in paraxial approximation
If the spread of the transverse moment $\Delta q$ is small compared to $p$, for $x>x_{0}$ we have
@@
now accounts for beam’s broadening. Neglecting the broadening yields Eq.(41).